# TempR1: Improving Temporal Understanding of MLLMs via Temporal-Aware Multi-Task Reinforcement Learning

### v2

@@ 4.5 Ablation Study
Ablation on Localization Reward for TAL. To evaluate the effectiveness of our Type-3 localization reward design, we conduct an ablation study on ActivityNet-v1.3 [activitynet] for TAL in Table 3. Removing the instance number reward leads to a noticeable performance drop (70.6 → 60.8 in mF1), confirming that explicitly penalizing instance count mismatches helps the model better estimate the number of actions. Replacing the DP-based matching with a naive sequential matching strategy severely degrades performance to 45.4 mF1, demonstrating the importance of optimal alignment between predicted and ground-truth intervals. Overall, the combination of both components yields the best result, achieving 70.6 mF1 on ActivityNet-v1.3.
Influence of RL Algorithm. We evaluate the impact of different RL algorithms on multi-task performance in Table 4. All three methods achieve strong results across most tasks, highlighting the robustness of multi-task reinforcement learning. Notably, DAPO [dapo] underperforms on the ActivityNet-v1.3 [activitynet] TAL task in terms of mF1. This is likely due to its token-level policy gradient loss, which biases optimization toward samples with more output instances and affects overall dataset performance. In contrast, GSPO [gspo] adopts sequence-level optimization, which better balances long and short samples and achieves superior results on multi-segment prediction tasks such as VHD and TAL. To maintain alignment with prior works [videochat-r1, timer1], we use GRPO [deepseekmath] as our default setting.
@@ 4.6 Qualitative Results
We present qualitative comparisons in Figure 4 to analyze the effectiveness of the proposed TempR1 model. In Figure 4(a), we compare two matching strategies used for computing the matching reward in the TAL task. The DP-based strategy produces more accurate and consistent temporal localization results in challenging multi-instance scenarios, demonstrating its ability to establish finer correspondences between predicted intervals and ground-truth instances. This leads to more reliable optimization targets during training and improved temporal precision. In Figure 4(b), we compare TempR1 with VideoChat-R1 [videochat-r1] on NExTGQA [nextgqa]. In the first example, both models predict the correct answer, but TempR1 provides more complete temporal evidence. In the second example, TempR1 shows improved temporal grounding and better reasoning consistency between the visual evidence and textual answer.

### v3

@@ 4.5 Ablation Study
Ablation on Localization Reward for TAL. To evaluate the effectiveness of our Type-3 localization reward design, we conduct an ablation study on ActivityNet-v1.3 [6] for TAL in Table 3. Removing the instance number reward leads to a noticeable performance drop (70.6 → 60.8 in mF1), confirming that explicitly penalizing instance count mismatches helps the model better estimate the number of actions. Replacing the DP-based matching with a naive sequential matching strategy severely degrades performance to 45.4 mF1, demonstrating the importance of optimal alignment between predicted and ground-truth intervals. Overall, the combination of both components yields the best result, achieving 70.6 mF1 on ActivityNet-v1.3.
Influence of RL Algorithm. We evaluate the impact of different RL algorithms on multi-task performance in Table 4. All three methods achieve strong results across most tasks, highlighting the robustness of multi-task reinforcement learning. Notably, DAPO [66] underperforms on the ActivityNet-v1.3 [6] TAL task in terms of mF1. This is likely due to its token-level policy gradient loss, which biases optimization toward samples with more output instances and affects overall dataset performance. In contrast, GSPO [72] adopts sequence-level optimization, which better balances long and short samples and achieves superior results on multi-segment prediction tasks such as VHD and TAL. To maintain alignment with prior works [27, 56], we use GRPO [45] as our default setting.
@@ 4.6 Qualitative Results
We present qualitative comparisons in Figure 4 to analyze the effectiveness of the proposed TempR1 model. In Figure 4(a), we compare two matching strategies used for computing the matching reward in the TAL task. The DP-based strategy produces more accurate and consistent temporal localization results in challenging multi-instance scenarios, demonstrating its ability to establish finer correspondences between predicted intervals and ground-truth instances. This leads to more reliable optimization targets during training and improved temporal precision. In Figure 4(b), we compare TempR1 with VideoChat-R1 [27] on NExTGQA [62]. In the first example, both models predict the correct answer, but TempR1 provides more complete temporal evidence. In the second example, TempR1 shows improved temporal grounding and better reasoning consistency between the visual evidence and textual answer.